OpenStudy (anonymous):

A system of equations is given as follows: x+y+2z=1 kx+2y+4z=k For what values of k does the system have an infinite number of solutions? Determine the solution to the system for this value of k.

OpenStudy (psymon):

Well, infinite solutions means you do elimination or substiution and you end up with 2 sides being perfectly equal. It means they'll be the same line essentially. So we want values of k that can either make everything cancel or make both sides equal after we do elimination. So it looks like, based on the y and z values, that our k line may just be a multiple of 2 of the first line. So if we let k = 2, we get: x + y + 2z = 1 2x + 2y + 4z = 2 Now this looks perfect. By elimination method, multiply the top equation by -2 and eliminate everything :3

OpenStudy (anonymous):

yes! I'm right(^_^) thanks, just confirming if I understand this correctly..

OpenStudy (psymon):

Yeah, np :P Did we leave our vectors section? xD

OpenStudy (anonymous):

nope, this is still vectors, planes and etc.

OpenStudy (psymon):

Ah, okay xD Yeah, Im still stuck there....because im going slow, haha. Almost on to the next section in the chapter -_-

OpenStudy (anonymous):

good for you!☺, i'm waiting for you a while ago, to answer my questions:)

OpenStudy (psymon):

Well, whatever you got that needs answering :3

OpenStudy (anonymous):

Determine the vector equation of the line that passes through A(-2, 3,6) and is parallel to the line of intersection of the planes \[\pi _{1}:2x-y+z=0\] and \[\pi _{2}: y+4z=0\]

OpenStudy (psymon):

Well, the line of intersection of theplanes requires a system of equations using the two plane equations themselves. From there we need tosolve for x, y, and z and put it in terms of t so we can have the line in parametricform.

OpenStudy (anonymous):

brb, i need to do something..

OpenStudy (psymon):

kkz

OpenStudy (anonymous):

, i'm back:) i'm doing another question and i'm stuck this is the question: determine the solution of this system of equation: 2x-y+2z=2 and -x+2y+z=1

OpenStudy (psymon):

All it says is solution, it doesnt say line of intersection or anything?

OpenStudy (anonymous):

nope..

OpenStudy (psymon):

Yeah, often times you find a line of intersection solution. Alright, lemme take a look then.

OpenStudy (anonymous):

sure:)

OpenStudy (psymon):

Yeah, I can get ya a line of intersection in parametric form, but if thats what is needed, doubt it, lol.

OpenStudy (anonymous):

that's fine.. let me see:)

OpenStudy (anonymous):

looks like it is a very long solution XD

OpenStudy (psymon):

So this is kind of a way you would go about finding a line of intersection. The first thing we do is eliminate one of the 3 variables and then choose to solve for one of the remaining two variables. It is not necessary that we come up with a numerical solution, just need to solve for a variable. Once we solve for a variable, we substitute this into one of the two original plane equations and solve for one of the other variables. The second variable we solve for, say we get x = z/7, we let that z/7 value = t. Then we solve for all 3 variables in terms of t. You'll understand the process as we go, though. 2x - y + 2z = 2 -x + 2y + z = -1 So we can just eliminate one of the variables to start with. I'll just eliminate y by multiplying the top plane by 2 4x - 2y + 4z = 4 -x + 2y + z = -1 = 3x + 5z = 3 So now we can just solve for either variable. I'll just solve for x saying: \[x=\frac{ 3-5z }{ 3 }\] So this x value is substituted into one of the two plane equations. Since the bottom equation only has a negative x I can do that one: \[-\frac{ 3-5z }{ 3 }+2y+z=-1\] = - 3 + 5z + 6y + 3z = -3 8z + 6y = 0 \[z=\frac{ -3y }{ 4 }\] Now we say that t = -3y/4 Therefore z = t y = -4t/3 x= 1- 5t/3 Kinda messy ish, but thats kinda the method xDD

OpenStudy (anonymous):

Oh WOW!=O give me time to read this

OpenStudy (psymon):

Yeah, np xD

OpenStudy (anonymous):

aah, I found my mistake, instead of multiplying first equation by 2, I multiply -2 to the second one:) thanks again..

OpenStudy (anonymous):

i mean 2.. to eliminate x

OpenStudy (psymon):

Lol, easy to make mistakes for sure xD

OpenStudy (anonymous):

i wish i could be like you:)

OpenStudy (psymon):

Lol, got many people on here way more advanced in math than I am, though, lol.

OpenStudy (anonymous):

at least your very helpful to me,XD

OpenStudy (psymon):

I try xD Just got a lot to learn, lol.

OpenStudy (anonymous):

lol, I peek to the next section, IT'S THREE PLANES!, it looks very hard O_0, did you do three planes?

OpenStudy (psymon):

Nope. Ive been going too slow. Ive been doing, like, all the rpoblems in this one section because I need to get it down xD

OpenStudy (anonymous):

same here:)

OpenStudy (anonymous):

wait, i checked the answer, it's different.. x=5/4 s y=s z=1-3/4 s

OpenStudy (psymon):

They solved for different variables. Its the same thing, but they have y as the main instead of z like we had.

OpenStudy (anonymous):

ok?, i'll check again

OpenStudy (psymon):

Can just solve for different variables and see what happens. Im not sure if theres a requirement or specific way to say solve for THIS ONE

OpenStudy (anonymous):

i got \[x=\frac{ 5-5z }{ 3}\]?